OpenStudy (anonymous):

Ok, i need a second opinion: find the maximum y value on the graph of y=f(x) f(x)=-x^2+8x+10. my answer is y=58....am i right?

OpenStudy (anonymous):

f'(x) = -2x + 8 0 = -2x + 8 x = 4 so f(4) = -(4)^2 + 8(4) + 10 = 58

OpenStudy (anonymous):

sweet, so i was right?

OpenStudy (anonymous):

yase

OpenStudy (noelgreco):

Oops. -4^2 = -16 The max is 26.